# Why Reasoning Matters? A Survey of Advancements in Multimodal Reasoning (v1)

### v1

@@ 4.4 Feedback
VOLCANO employs a three-stage iterative framework—initial response generation, visual self-assessment, and revision—to reduce hallucinations by emphasizing accurate image details (Lee et al., 2024). Similarly, LOOKBACK mandates explicit atomic verification against images to enhance both critique and correction processes for LVLMs (Wu et al., 2024b). LLaVA-Critic advances feedback by incorporating diverse critic instructions and iterative DPO training, leveraging internal data (Xiong et al., 2024). Additionally, tailored visual prompting with iterative binary verification effectively enhances semantic grounding in models like LLaVA-1.5, ViP-LLaVA, and CogVLM (Liao et al., 2024a). Minimalist reinforcement learning setups have shown effectiveness, particularly for multimodal mathematical reasoning, with larger models like InternVL2.5-38B benefiting from difficulty-based data filtering to stabilize training (Meng et al., 2025). Explicit visual CoT mechanisms, exemplified by VisCoT, integrate bounding box predictions and iterative cropping to enhance visual question-answering performance (Shao et al., 2024). In video understanding, VIDEOTREE utilizes adaptive hierarchical clustering for efficient navigation through structured tree-based representations (Wang et al., 2024i). Lastly, prompt engineering strategies like MiHO and MiHI significantly reduce hallucinations without model retraining, post-processing evaluations with GPT-4 further enhancing reliability without architectural modifications (Han et al., 2024; Hu et al., 2025).
@@ 4.5 Iterative Refinement
Adaptive inference transforms vision-language reasoning by facilitating dynamic, context-sensitive processing to enhance accuracy and efficiency. Central to this approach is iterative output refinement through internal feedback and external verification. For instance, IXC-2.5 employs Best-of-N sampling guided by reward models to optimize performance and response length (Zang et al., 2025), while PARM++ uses reflection-based refinement for better alignment between images and prompts (Guo et al., 2025). Further advancements include LLaVA-o1’s structured perception and stage-level beam search for decomposing complex reasoning tasks (Xu et al., 2025). To mitigate multi-modal hallucinations, MEMVR implements visual retracing based on uncertainty to optimize the cost-accuracy tradeoff (Zou et al., 2024a), MVP aggregates certainty across diverse views (Qu et al., 2024), and SID employs token-level contrastive decoding to eliminate irrelevant content (Huo et al., 2024). Moreover, DualFocus integrates macro- and micro-level reasoning for precise attention control (Cao et al., 2024b), LISA++ enables test-time compute scaling without retraining (Yang et al., 2024a), and PerceptionGPT accelerates inference through adaptive token embeddings encoding spatial cues and dynamically weighted layers (Pi et al., 2023).
@@ 4.6 Dataset Curation
Recent work shows that high-quality, strategically curated datasets outperform large but noisy alternatives in guiding reasoning paths. R-CoT (Deng et al., 2024) enhances geometric reasoning via reverse generation and stepwise synthesis. Task-specific datasets such as SMIR (Li et al., 2025a) show up to 8% performance gains over generic datasets, demonstrating the value of targeted curation. Gradual complexity in dataset construction, as shown in VIREO (Cheng et al., 2024a), ensures foundational reasoning skills before introducing advanced tasks. Scalable dataset creation through automation also proves effective. VideoVista (Li et al., 2024e) uses auto-annotations for video reasoning, outperforming manually curated sets. DecoVQA+ (Zhang et al., 2024b) explicitly teaches when to apply question decomposition, aided by SelectiveVQD loss for better strategy selection. LocVLM (Ranasinghe et al., 2024) scales with pseudo-data and implicit feedback signals like location and relevance prediction. CogCoM (Qi et al., 2024) incorporates tasks like grounding and manipulation to reinforce step-by-step reasoning.
@@ 5 Datasets and Benchmarks
Structured and Task-Specific Reasoning Datasets such as Visual-RFT(Liu et al., 2025c), CapQA(Hu et al., 2025), GUIDE(Liang et al., 2024a), STAR(Wu et al., 2024a), Visual Genome(Zhang et al., 2024f), VL-GPT(Zhu et al., 2023), and Interfacing (Zou et al., 2024b) have introduced structured prompting that significantly enhances models’ reasoning accuracy by explicitly guiding reasoning processes. In contrast to earlier benchmarks like STAR, these datasets place a stronger emphasis on task-specific reward functions and structured inference, aiming to improve generalization across diverse visual and linguistic contexts.
@@
Temporal and Spatial Reasoning Temporal benchmarks like VisualQA, and TemporalVQA (Imam et al., 2025), TLQA (Swetha et al., 2025), REXTIME (Chen et al., 2024c), FrameCap (Xue et al., 2024), and VideoVista (Li et al., 2024e) have revealed substantial limitations in current multimodal language models, achieving accuracy significantly below human levels (15% versus 90%). These benchmarks emphasize the necessity for temporal logic annotations and highlight the difficulty models face when reasoning across video frames. In parallel, spatial datasets such as SpatialVLM (Chen et al., 2024a), WhatsUp (Kamath et al., 2023), DC2 (Wang et al., 2024e), and Grounded (Chen et al., 2024f) showcase improvements through extensive spatial reasoning QA pairs and challenging spatial configurations.
@@
Iterative and Reflective Reasoning Iterative reasoning capabilities have markedly improved through datasets like VISCO (Wu et al., 2024b), Mulberry-260k (Yao et al., 2024), FIRE (Li et al., 2024d), VLFeedback (Li et al., 2024c), Silkie (Li et al., 2023b), TIIL (Huang et al., 2024c), ConMe (Huang et al., 2024a), and Reflective (Zhang et al., 2024f), enabling models to learn from mistakes by both incorporating reflective and iterative feedback loops. For instance, VISCO demonstrated that human critiques significantly outperform model critiques (76% error correction), highlighting a persistent gap in models’ ability to independently self-correct.
@@
Complex Multimodal Evaluation Complex multimodal datasets like EMMA (Hao et al., 2025), CoMT (Cheng et al., 2024c), SMIR (Li et al., 2025a), ProVision (Zhang et al., 2024e), MAGEBench (Zhang et al., 2024i), MM-Vet v2 (Yu et al., 2024), JourneyBench (Wang et al., 2025), VERIFY (Bi et al., 2025) and CompCap (Chen et al., 2024g) exposed limitations in current multimodal models’ integration capabilities. EMMA specifically shows that models struggle significantly with iterative multimodal interactions, particularly when tasks demand deep integration across domain-specific problems. Benchmarks like FineCops-Ref (Liu et al., 2025b), VL-RewardBench (Li et al., 2024b), MM-SAP (Wang et al., 2024h), PCA-Bench (Chen et al., 2024d), AttCoSeg (Pramanick et al., 2024), and M3CoT (Chen et al., 2024e) emphasize detailed evaluation metrics (Recall@k, AUROC). These benchmarks provide insights into models’ actual reasoning capabilities beyond traditional accuracy metrics.
@@
Counterfactual and Logical Reasoning Datasets like CounterCurate (Zhang et al., 2024c), C-VQA (Zhang et al., 2024g), CausalChaos! (Parmar et al., 2024), LogicAI (Xiao et al., 2024), and MCDGRAPH (Zhu et al., 2024) explicitly test model reasoning robustness, showing improvements through challenging counterfactual examples and structured logical questions.
@@
Active Perception and Progressive Reasoning Active perception datasets such as ActiView (Wang et al., 2024j) and progressive reasoning benchmarks such as Blink (Fu et al., 2024b), ADL-X (Reilly et al., 2024), and GUIDE (Liang et al., 2024a) challenge models’ dynamic reasoning capabilities through active view adjustments and frame-level progression.
@@ Appendix A Implications from the Learning Science Perspective
One key theory is Cognitive Load Theory (Sweller, 2011; Paas et al., 2004), which reduces a human’s working memory workload by breaking complex problems into manageable steps. CoT reasoning follows this approach by prompting models to generate step-by-step solutions rather than arriving at an answer in one leap. This structured decomposition mirrors how human learners handle intricate problems by offloading cognitive effort across multiple processing stages. Similarly, in the context of learning, scaffolding refers to offering temporary assistance to learners, enabling them to accomplish tasks that would be difficult or impossible for them to complete independently. This support is gradually removed as learners develop the skills and confidence required to perform the task on their own (Van de Pol et al., 2010). This supports CoT’s design by guiding LLMs through intermediate steps, much like how educators provide structured support to help students grasp new concepts before transitioning to independent problem-solving. These theories are both rooted in the Zone of Proximal Development (ZPD), a key construct in Lev Vygotsky’s theory of learning and development (Vygotsky, 1978).
CoT reasoning also aligns with metacognition (Lai, 2011) and self-explanation (Bisra et al., 2018; Chi et al., 1994) by encouraging models to “think about their thinking”. Just as self-explanation improves human learning by prompting individuals to articulate their reasoning (Chi et al., 1994; VanLehn et al., 1992), CoT forces LLMs to justify their steps, reducing reliance on shallow heuristics. This ties into Dual-Process Theory (Frankish, 2010), where CoT shifts LLMs from fast, intuitive decision-making (System 1) to deliberate, analytical reasoning (System 2), leading to more logical and consistent outputs.
Moreover, CoT fosters constructivist learning (Waite-Stupiansky, 2022; Bruner, 2009) by enabling LLMs to incrementally build knowledge structures through active reasoning. Instead of passively retrieving answers from training data, the model synthesizes prior knowledge with new information, improving adaptability. Additionally, analogical reasoning (Gentner & Maravilla, 2017) plays a role in CoT by helping LLMs map relationships between concepts, allowing them to generalize problem-solving strategies across different contexts.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=10em,font=,, where level=3text width=8.2em,font=,, [ Inference [ Adaptive
Inference [  Guo et al. (2025),  Hu et al. (2025),  Yang et al. (2024b),
 Xue et al. (2024),  Wu et al. (2024b),  Wang et al. (2024f),
 Zou et al. (2024a),  Qu et al. (2024),  Zhang et al. (2024k),
 Huo et al. (2024),  Liao et al. (2024b),  Su et al. (2024),
 Li et al. (2024e),  Wang et al. (2024e),  Xu et al. (2025),
 Zhao et al. (2024),  Liao et al. (2024a),  Mahmood et al. (2024),
 Cao et al. (2024b),  Wu et al. (2024a),  Yang et al. (2024a),
 Li et al. (2023a),  Pi et al. (2023) , leaf, text width=30em ] ] [ Reward
Models [  Guo et al. (2025),  Hao et al. (2025),  Li et al. (2024b),
 Fu et al. (2024c),  Ke et al. (2024),  Xiong et al. (2024),
 Cascante-Bonilla et al. (2024),  Cheng et al. (2024b),
 Zhang et al. (2024k), , leaf, text width=30em ] ] [ Feedback [  Guo et al. (2025),  Wu et al. (2024b),  Hu et al. (2025),
 Meng et al. (2025),  Li et al. (2024d),  Xiong et al. (2024),
 Cheng et al. (2024b),  Zhang et al. (2024k),  Zheng et al. (2024),
 Liao et al. (2024a),  Han et al. (2024),  Wang et al. (2024i),
 Shao et al. (2024),  Liang et al. (2024b),
 Lee et al. (2024),  Hu et al. (2024), , leaf, text width=30em ] ] [ Multimodal
Decomposition [  Xiang et al. (2025),  Liang et al. (2025),  Liu et al. (2025a),
 Shlapentokh-Rothman et al. (2024),  Zhang et al. (2024b),
 Chen et al. (2024f),  Hsu et al. (2024),  Su et al. (2024),
 Reilly et al. (2024),  Wang et al. (2024a),  Zhang et al. (2024h),
 Min et al. (2024),  Parmar et al. (2024),
 Wang et al. (2024b),  Dzabraev et al. (2024),  Wu et al. (2024a),
 Wu & Xie (2023),  Zhang et al. (2023), , leaf, text width=30em ] ] [ Efficiency and
Scalability [  Goyal et al. (2025),  Li et al. (2025b),  Sun et al. (2024),
 Fan et al. (2024),  Park et al. (2025),  Wang et al. (2024e),
 Wang et al. (2025),  Huo et al. (2024),
 Lin et al. (2024),  Ni et al. (2024),  Wang et al. (2024i),
 Huang et al. (2024a),  Zhang et al. (2024m),  Zhan et al. (2024),
 Chen et al. (2024b),  Wang et al. (2024b),  Zhao et al. (2023),
 Pi et al. (2023),  Chen et al. (2024a) , leaf, text width=30em ] ] [ Search
Strategies [  Yao et al. (2024),  Yang et al. (2024b),  Wang et al. (2024f),
 Imam et al. (2025),  Xu et al. (2025),  Qu et al. (2024),
 Wang et al. (2024e),  Li et al. (2024e),  Kazemi et al. (2024),
 Shin et al. (2024),  Zhang et al. (2024l),
 Zhao et al. (2024),  Fu et al. (2024c),  Mahmood et al. (2024),
 Wu & Xie (2023),  Hu et al. (2023), , leaf, text width=30em ] ] ]
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=10em,font=,, where level=3text width=8.2em,font=,, [ Post
Training [ Policy
Optimization [  Fu et al. (2024c),  Zhang et al. (2024d),  Hu et al. (2024),
 Tian et al. (2024),  Kim et al. (2023),  Yan et al. (2024),
 Ouali et al. (2024),  Xiong et al. (2024),  Wang et al. (2024j),
 Waite et al. (2025),  Wei et al. (2025),  Zhao et al. (2025),
 Xu et al. (2025),  Deng et al. (2024), , leaf, text width=30em ] ] [ Model
Architecture [  Reilly et al. (2024),  Ma et al. (2024a),  Zhang et al. (2024j),
 Zhan et al. (2024),  Jiao et al. (2024),
 He et al. (2024),  Ma et al. (2023),  Ben-Kish et al. (2024),
 Hu et al. (2024),  Wang et al. (2024d),  Wang et al. (2024c),
 Cao et al. (2024a),  Chen et al. (2024f),  Huang et al. (2024d), , leaf, text width=30em ] ] [ Reward
Alignment [  Hao et al. (2025),  Li et al. (2024b),  Xiong et al. (2024),
 Ke et al. (2024),  Fu et al. (2024c),  Dzabraev et al. (2024),
 Hu et al. (2024), , leaf, text width=30em ] ] [ Iterative
Refinement [  Meng et al. (2025),  Zheng et al. (2024),  Liao et al. (2024a),
 Wu et al. (2024b),  Shao et al. (2024),
 Han et al. (2024),  Hu et al. (2024),  Lee et al. (2024),
, leaf, text width=30em ] ] [ Spatial
Temporal
Modeling [  Wang et al. (2024i),  Zhang et al. (2024m),  Chen et al. (2024a),
 Wu & Xie (2023), , leaf, text width=30em ] ] [ Dataset
Curation [  Fan et al. (2024),  Wang et al. (2025),
 Nagarajan & Torresani (2024),  Zhang et al. (2024l),
 Lin et al. (2024),  Zhao et al. (2023),  Hu et al. (2023), , leaf, text width=30em ] ] ]

### v2

@@ 4.4 Feedback
VOLCANO employs a three-stage iterative framework—initial response generation, visual self-assessment, and revision—to reduce hallucinations by emphasizing accurate image details (2311.07362-lee_volcano:_2024). Similarly, LOOKBACK mandates explicit atomic verification against images to enhance both critique and correction processes for LVLMs (2412.02172-wu_visco:_2024). LLaVA-Critic advances feedback by incorporating diverse critic instructions and iterative DPO training, leveraging internal data (2410.02712-xiong_llava-critic:_2024). Additionally, tailored visual prompting with iterative binary verification effectively enhances semantic grounding in models like LLaVA-1.5, ViP-LLaVA, and CogVLM (2404.06510-liao_can_2024). Minimalist reinforcement learning setups have shown effectiveness, particularly for multimodal mathematical reasoning, with larger models like InternVL2.5-38B benefiting from difficulty-based data filtering to stabilize training (2503.07365-meng_mm-eureka:_2025). Explicit visual CoT mechanisms, exemplified by VisCoT, integrate bounding box predictions and iterative cropping to enhance visual question-answering performance (2403.16999-shao_visual_2024). In video understanding, VIDEOTREE utilizes adaptive hierarchical clustering for efficient navigation through structured tree-based representations (2405.19209-wang_videotree:_2024). Lastly, prompt engineering strategies like MiHO and MiHI significantly reduce hallucinations without model retraining, post-processing evaluations with GPT-4 further enhancing reliability without architectural modifications (2402.01345-han_skip_2024; 2501.02964-hu_socratic_2025).
@@ 4.5 Iterative Refinement
Adaptive inference transforms vision-language reasoning by facilitating dynamic, context-sensitive processing to enhance accuracy and efficiency. Central to this approach is iterative output refinement through internal feedback and external verification. For instance, IXC-2.5 employs Best-of-N sampling guided by reward models to optimize performance and response length (2501.12368-zang_internlm-xcomposer2.5-reward:_2025), while PARM++ uses reflection-based refinement for better alignment between images and prompts (2501.13926-guo_can_2025). Further advancements include LLaVA-o1’s structured perception and stage-level beam search for decomposing complex reasoning tasks (2411.10440-xu_llava-cot:_2025). To mitigate multi-modal hallucinations, MEMVR implements visual retracing based on uncertainty to optimize the cost-accuracy tradeoff (2410.03577-zou_look_2024), MVP aggregates certainty across diverse views (2408.17150-qu_look_2024), and SID employs token-level contrastive decoding to eliminate irrelevant content (2408.02032-huo_self-introspective_2024). Moreover, DualFocus integrates macro- and micro-level reasoning for precise attention control (2402.14767-cao_dualfocus:_2024), LISA++ enables test-time compute scaling without retraining (2312.17240-yang_lisa++:_2024), and PerceptionGPT accelerates inference through adaptive token embeddings encoding spatial cues and dynamically weighted layers (2311.06612-pi_perceptiongpt:_2023).
@@ 4.6 Dataset Curation
Recent work shows that high-quality, strategically curated datasets outperform large but noisy alternatives in guiding reasoning paths. R-CoT (2410.17885-deng_r-cot:_2024) enhances geometric reasoning via reverse generation and stepwise synthesis. Task-specific datasets such as SMIR (2501.03675-li_smir:_2025) show up to 8% performance gains over generic datasets, demonstrating the value of targeted curation. Gradual complexity in dataset construction, as shown in VIREO (2406.19934-cheng_least_2024), ensures foundational reasoning skills before introducing advanced tasks. Scalable dataset creation through automation also proves effective. VideoVista (2406.11303-li_videovista:_2024) uses auto-annotations for video reasoning, outperforming manually curated sets. DecoVQA+ (2409.19339-zhang_visual_2024) explicitly teaches when to apply question decomposition, aided by SelectiveVQD loss for better strategy selection. LocVLM (2404.07449-ranasinghe_learning_2024) scales with pseudo-data and implicit feedback signals like location and relevance prediction. CogCoM (2402.04236-qi_cogcom:_2024) incorporates tasks like grounding and manipulation to reinforce step-by-step reasoning.
@@ 5 Datasets and Benchmarks
Structured and Task-Specific Reasoning Datasets such as Visual-RFT(2503.01785-liu_visual-rft:_2025), CapQA(2501.02964-hu_socratic_2025), GUIDE(2406.18227-liang_guide:_2024), STAR(2405.09711-wu_star:_2024), Visual Genome(2407.11422-zhang_reflective_2024), VL-GPT(2312.09251-zhu_vl-gpt:_2023), and Interfacing (2312.07532-zou_interfacing_2024) have introduced structured prompting that significantly enhances models’ reasoning accuracy by explicitly guiding reasoning processes. In contrast to earlier benchmarks like STAR, these datasets place a stronger emphasis on task-specific reward functions and structured inference, aiming to improve generalization across diverse visual and linguistic contexts.
@@
Temporal and Spatial Reasoning Temporal benchmarks like VisualQA, and TemporalVQA (2501.10674-imam_can_2025), TLQA (2501.07214-swetha_timelogic:_2025), REXTIME (2406.19392-chen_rextime:_2024), FrameCap (2412.02071-xue_progress-aware_2024), and VideoVista (2406.11303-li_videovista:_2024) have revealed substantial limitations in current multimodal language models, achieving accuracy significantly below human levels (15% versus 90%). These benchmarks emphasize the necessity for temporal logic annotations and highlight the difficulty models face when reasoning across video frames. In parallel, spatial datasets such as SpatialVLM (2401.12168-chen_spatialvlm:_2024), WhatsUp (2310.19785-kamath_whats_2023), DC2 (2408.15556-wang_divide_2024), and Grounded (2408.14469-chen_grounded_2024) showcase improvements through extensive spatial reasoning QA pairs and challenging spatial configurations.
@@
Iterative and Reflective Reasoning Iterative reasoning capabilities have markedly improved through datasets like VISCO (2412.02172-wu_visco:_2024), Mulberry-260k (2412.18319-yao_mulberry:_2024), FIRE (2407.11522-li_fire:_2024), VLFeedback (2410.09421-li_vlfeedback:_2024), Silkie (2312.10665-li_silkie:_2023), TIIL (2404.18033-huang_exposing_2024), ConMe (2406.08164-huang_conme:_2024), and Reflective (2407.11422-zhang_reflective_2024), enabling models to learn from mistakes by both incorporating reflective and iterative feedback loops. For instance, VISCO demonstrated that human critiques significantly outperform model critiques (76% error correction), highlighting a persistent gap in models’ ability to independently self-correct.
@@
Complex Multimodal Evaluation Complex multimodal datasets like EMMA (2501.05444-hao_can_2025), CoMT (2412.12932-cheng_comt:_2024), SMIR (2501.03675-li_smir:_2025), ProVision (2412.07012-zhang_provision:_2024), MAGEBench (2412.04531-zhang_magebench:_2024), MM-Vet v2 (2408.00765-yu_mm-vet_2024), JourneyBench (2409.12953-wang_journeybench:_2025), VERIFY (bi2025verify) and CompCap (2412.05243-chen_compcap:_2024) exposed limitations in current multimodal models’ integration capabilities. EMMA specifically shows that models struggle significantly with iterative multimodal interactions, particularly when tasks demand deep integration across domain-specific problems. Benchmarks like FineCops-Ref (2409.14750-liu_finecops-ref:_2025), VL-RewardBench (2411.17451-li_vlrewardbench:_2024), MM-SAP (2401.07529-wang_mm-sap:_2024), PCA-Bench (2402.15527-chen_pca-bench:_2024), AttCoSeg (2312.12423-pramanick_jack_2024), and M3CoT (2405.16473-chen_m\$^3\$cot:_2024) emphasize detailed evaluation metrics (Recall@k, AUROC). These benchmarks provide insights into models’ actual reasoning capabilities beyond traditional accuracy metrics.
@@
Counterfactual and Logical Reasoning Datasets like CounterCurate (2402.13254-zhang_countercurate:_2024), C-VQA (2310.06627-zhang_what_2024), CausalChaos! (2404.01299-parmar_causalchaos!_2024), LogicAI (2407.04973-xiao_logicvista:_2024), and MCDGRAPH (2412.13540-zhu_benchmarking_2024) explicitly test model reasoning robustness, showing improvements through challenging counterfactual examples and structured logical questions.
@@
Active Perception and Progressive Reasoning Active perception datasets such as ActiView (2410.04659-wang_actiview:_2024) and progressive reasoning benchmarks such as Blink (2404.12390-fu_blink:_2024), ADL-X (2406.09390-reilly_llavidal:_2024), and GUIDE (2406.18227-liang_guide:_2024) challenge models’ dynamic reasoning capabilities through active view adjustments and frame-level progression.
@@ Appendix A Implications from the Learning Science Perspective
One key theory is Cognitive Load Theory (sweller2011cognitive; paas2004cognitive), which reduces a human’s working memory workload by breaking complex problems into manageable steps. CoT reasoning follows this approach by prompting models to generate step-by-step solutions rather than arriving at an answer in one leap. This structured decomposition mirrors how human learners handle intricate problems by offloading cognitive effort across multiple processing stages. Similarly, in the context of learning, scaffolding refers to offering temporary assistance to learners, enabling them to accomplish tasks that would be difficult or impossible for them to complete independently. This support is gradually removed as learners develop the skills and confidence required to perform the task on their own (van2010scaffolding). This supports CoT’s design by guiding LLMs through intermediate steps, much like how educators provide structured support to help students grasp new concepts before transitioning to independent problem-solving. These theories are both rooted in the Zone of Proximal Development (ZPD), a key construct in Lev Vygotsky’s theory of learning and development (vygotsky1978mind).
CoT reasoning also aligns with metacognition (lai2011metacognition) and self-explanation (bisra2018inducing; chi1994eliciting) by encouraging models to “think about their thinking”. Just as self-explanation improves human learning by prompting individuals to articulate their reasoning (chi1994eliciting; vanlehn1992model), CoT forces LLMs to justify their steps, reducing reliance on shallow heuristics. This ties into Dual-Process Theory (frankish2010dual), where CoT shifts LLMs from fast, intuitive decision-making (System 1) to deliberate, analytical reasoning (System 2), leading to more logical and consistent outputs.
Moreover, CoT fosters constructivist learning (waite2022jean; bruner2009process) by enabling LLMs to incrementally build knowledge structures through active reasoning. Instead of passively retrieving answers from training data, the model synthesizes prior knowledge with new information, improving adaptability. Additionally, analogical reasoning (gentner2017analogical) plays a role in CoT by helping LLMs map relationships between concepts, allowing them to generalize problem-solving strategies across different contexts.
@@
Inference
Adaptive Inference 2501.13926-guo_can_2025,  2501.02964-hu_socratic_2025,  2412.10471-yang_vca:_2024, 2412.02071-xue_progress-aware_2024,  2412.02172-wu_visco:_2024,  2412.03704-wang_scaling_2024, 2410.03577-zou_look_2024,  2408.17150-qu_look_2024,  2410.16198-zhang_improve_2024, 2408.02032-huo_self-introspective_2024,  2407.03008-liao_align_2024,  2406.10923-su_investigating_2024, 2406.11303-li_videovista:_2024,  2408.15556-wang_divide_2024,  2411.10440-xu_llava-cot:_2025, 2405.16071-zhao_dynrefer:_2024,  2404.06510-liao_can_2024,  2403.14743-mahmood_vurf:_2024, 2402.14767-cao_dualfocus:_2024,  2405.09711-wu_star:_2024,  2312.17240-yang_lisa++:_2024, 2311.13601-li_visual_2023,  2311.06612-pi_perceptiongpt:_2023 Reward Models 2501.13926-guo_can_2025,  2501.05444-hao_can_2025,  2411.17451-li_vlrewardbench:_2024, 2406.00645-fu_furl:_2024,  2403.12884-ke_hydra:_2024,  2410.02712-xiong_llava-critic:_2024, 2410.22315-cascante-bonilla_natural_2024,  2411.00855-cheng_vision-language_2024, 2410.16198-zhang_improve_2024, Feedback 2501.13926-guo_can_2025,  2412.02172-wu_visco:_2024,  2501.02964-hu_socratic_2025, 2503.07365-meng_mm-eureka:_2025,  2407.11522-li_fire:_2024,  2410.02712-xiong_llava-critic:_2024, 2411.00855-cheng_vision-language_2024,  2410.16198-zhang_improve_2024,  2411.12591-zheng_thinking_2024, 2404.06510-liao_can_2024,  2402.01345-han_skip_2024,  2405.19209-wang_videotree:_2024, 2403.16999-shao_visual_2024,  2404.04007-liang_neural-symbolic_2024, 2311.07362-lee_volcano:_2024,  2312.03052-hu_visual_2024, Multimodal Decomposition 2503.06252-xiang_can_2025,  2501.13536-liang_reasvqa:_2025,  2501.05069-liu_commonsense_2025, 2412.08564-shlapentokh-rothman_can_2024,  2409.19339-zhang_visual_2024, 2408.14469-chen_grounded_2024,  2409.08202-hsu_what_2024,  2406.10923-su_investigating_2024, 2406.09390-reilly_llavidal:_2024,  2408.05019-wang_instruction_2024,  2406.16620-zhang_omagent:_2024, 2404.06511-min_morevqa:_2024,  2404.01299-parmar_causalchaos!_2024, 2403.09394-wang_git:_2024,  2404.01911-dzabraev_vlrm:_2024,  2405.09711-wu_star:_2024, 2312.14135-wu_v*:_2023,  2309.15112-zhang_internlm-xcomposer:_2023, Efficiency and Scalability 2502.00382-goyal_masked_2025,  2501.07542-li_imagine_2025,  2412.08635-sun_multimodal_2024, 2412.18072-fan_mmfactory:_2024,  2501.02669-park_generalizing_2025,  2408.15556-wang_divide_2024, 2409.12953-wang_journeybench:_2025,  2408.02032-huo_self-introspective_2024, 2407.04681-lin_rethinking_2024,  2410.03321-ni_visual-o1:_2024,  2405.19209-wang_videotree:_2024, 2406.08164-huang_conme:_2024,  2403.07487-zhang_motion_2024,  2403.09333-zhan_griffon_2024, 2403.19322-chen_plug-and-play_2024,  2403.09394-wang_git:_2024,  2312.11420-zhao_tuning_2023, 2311.06612-pi_perceptiongpt:_2023,  2401.12168-chen_spatialvlm:_2024 Search Strategies 2412.18319-yao_mulberry:_2024,  2412.10471-yang_vca:_2024,  2412.03704-wang_scaling_2024, 2501.10674-imam_can_2025,  2411.10440-xu_llava-cot:_2025,  2408.17150-qu_look_2024, 2408.15556-wang_divide_2024,  2406.11303-li_videovista:_2024,  2406.09175-kazemi_remi:_2024, 2404.15190-shin_socratic_2024,  2405.03272-zhang_worldqa:_2024, 2405.16071-zhao_dynrefer:_2024,  2406.00645-fu_furl:_2024,  2403.14743-mahmood_vurf:_2024, 2312.14135-wu_v*:_2023,  2306.08129-hu_avis:_2023,
@@
Post
Training
Policy Optimization 2406.00645-fu_furl:_2024,  2402.07384-zhang_exploring_2024,  2312.03052-hu_visual_2024, 2310.07932-tian_what_2024,  2309.10790-kim_guide_2023,  2402.06118-yan_vigor:_2024, 2408.10433-ouali_clip-dpo:_2024,  2410.02712-xiong_llava-critic:_2024,  2410.04659-wang_actiview:_2024, 2501.18880-waite_rls3:_2025,  2503.08525-wei_gtr:_2025,  2503.05379-zhao_r1-omni:_2025, 2411.10440-xu_llava-cot:_2025,  2410.17885-deng_r-cot:_2024, Model Architecture 2406.09390-reilly_llavidal:_2024,  2404.13847-ma_eventlens:_2024,  2404.01258-zhang_direct_2024, 2403.09333-zhan_griffon_2024,  2401.17981-jiao_training-free_2024, 2401.03105-he_incorporating_2024,  2312.08870-ma_vista-llama:_2023,  2312.03631-ben-kish_mitigating_2024, 2312.03052-hu_visual_2024,  2311.03079-wang_cogvlm:_2024,  2410.09575-wang_reconstructive_2024, 2410.08209-cao_emerging_2024,  2408.14469-chen_grounded_2024,  2408.13890-huang_making_2024, Reward Alignment 2501.05444-hao_can_2025,  2411.17451-li_vlrewardbench:_2024,  2410.02712-xiong_llava-critic:_2024, 2403.12884-ke_hydra:_2024,  2406.00645-fu_furl:_2024,  2404.01911-dzabraev_vlrm:_2024, 2312.03052-hu_visual_2024, Iterative Refinement 2503.07365-meng_mm-eureka:_2025,  2411.12591-zheng_thinking_2024,  2404.06510-liao_can_2024, 2412.02172-wu_visco:_2024,  2403.16999-shao_visual_2024, 2402.01345-han_skip_2024,  2312.03052-hu_visual_2024,  2311.07362-lee_volcano:_2024, Spatial Temporal Modeling 2405.19209-wang_videotree:_2024,  2403.07487-zhang_motion_2024,  2401.12168-chen_spatialvlm:_2024, 2312.14135-wu_v*:_2023, Dataset Curation 2412.18072-fan_mmfactory:_2024,  2409.12953-wang_journeybench:_2025, 2404.16222-nagarajan_step_2024,  2405.03272-zhang_worldqa:_2024, 2407.04681-lin_rethinking_2024,  2312.11420-zhao_tuning_2023,  2306.08129-hu_avis:_2023,